### v1

# Properties of magnetic turbulence in GRB afterglows

Abstract
We present a model-independent way to characterise properties of the magnetic-field turbulence in the emitting regions of Gamma-Ray Burst afterglows. Our only assumption is that afterglows’ synchrotron radiation is efficient. It turns out that the gyroradius of plasma particles must be smaller (with a good margin) than the correlation length of the magnetic-field fluctuations. Such turbulence is essentially non-linear and therefore must be produced by some kind of MHD instability, likely acting on top of kinetic Weibel instability. We also find that the emitting particles are loosely confined to local magnetic-field structures and diffusion allows them to sample the entire distribution of local magnetization values. This means that one-zone approach to modelling the afterglow spectra is still valid despite the non-linear nature of the magnetic turbulence. However, the non-linear turbulence may (and likely will) change the synchrotron spectrum of individual electrons.
gamma-ray burst: general; radiation mechanisms: non-thermal; turbulence; magnetic fields
1 Introduction
Gamma-Ray Burst (GRB) afterglows are emission of relativistic shock waves formed as material ejected by the GRB’s central engine hits the circumburst medium. This complicated phenomenon could be broadly decomposed into three problems: the emission mechanism, the particle acceleration mechanism, and the shock’s microphysics.
Physics of the emission process itself is generally well understood — it finds explanation within the framework of the synchrotron-self-Compton (SSC) model (see, e.g. Sari et al., 1998; Sari & Esin, 2001; Nakar et al., 2009). The SSC model describes the bulk of afterglow spectrum (up to tens of GeV in photon energy) as synchrotron emission of energetic electrons (injected from some acceleration mechanism) in the magnetic field of the afterglow’s shock wave. Some of the synchrotron photons are upscattered by the same electrons to produce another, inverse Compton (IC) component, which extends to TeV energies. The available broad-band observations of GRB afterglows (MAGIC Collaboration et al., 2019; H. E. S. S. Collaboration et al., 2021) show that the synchrotron component dominates over the IC component. Hence, we do not include the latter in our analysis.
There are two options for the particle acceleration mechanism in relativistic shocks: a modification of common diffusive shock acceleration (DSA) (see, e.g. Achterberg et al., 2001; Keshet & Waxman, 2005) and the converter mechanism (Derishev et al., 2003), whose specific realisation for the SSC-emitting shocks is the pair-balance model (Derishev & Piran, 2016). In the pair-balance model the accelerated electrons (and positrons) are $e^{-}e^{+}$-pairs, produced in the upstream via collisions of IC photons with synchrotron photons, and then transported by the flow into the downstream gaining energy thereby. The pair-balance model is the most plausible candidate for the acceleration mechanism in GRB afterglows — unlike DSA, it correctly predicts the typical energy of the injected electrons (prediction about the spectral index are similar for both DSA and converter mechanisms).
The shock’s microphysics, and especially the process of magnetic field generation in initially weakly magnetized upstream fluid, is the least clear part of the puzzle. Yet it has implications for both the emission mechanism and the particle acceleration mechanism. Numeric simulations of collisionless and initially unmagnetized relativistic shocks successfully demonstrated generation of the magnetic field turbulence at the shock front, but this magnetic field is very short-lived compared to the synchrotron-loss timescale (see, e.g. Chang et al., 2008; Sironi et al., 2015). The pair-balance model may help resolving this issue — its distinctive feature is greatly extended in time injection of anisotropic particles into the upstream, that slows growth of the magnetic turbulence and potentially makes it persistent over the entire synchrotron cooling time. The possibility that slower growth of the magnetic turbulence leads to longer lifetime, suggested in the original paper, was later confirmed numerically (Garasev & Derishev, 2016). However, it remains unclear if a kinetic instability with the growth rate that matches the actual afterglow durations is capable of producing strong enough magnetic turbulence.
In this paper we present a way to characterize the magnetic turbulence in GRB afterglows using most general and model-independent approach. Our only assumption is that synchrotron radiation is efficient in GRB afterglows. Surprisingly, we are able to prove a rather strong statement the the magnetic turbulence in afterglow emitting regions is non-linear. This is unexpected and likely means that the present-day theories of shocks’ microphysics need a significant modification.
To keep our analysis more general, we treat the post-shock fluid as consisting of two components: the synchrotron-emitting electrons that produce the observed radiation, and the plasma particles (either ions or the emitting electrons themselves) that takes up most of the energy density and conveys currents supporting the magnetic field. Such a system is characterized by 6 parameters: plasma frequency $\omega_{\mathrm{pl}}$ (or, equivalently, plasma skin depth $\ell_{s}=c/\omega_{\mathrm{pl}}$), correlation length of the magnetic field fluctuations $\lambda_{c}$, decay time of the turbulent magnetic field $t_{\mathrm{dec}}$, gyroradius of plasma particles $r_{\mathrm{g,pl}}$, radiative energy loss timescale for emitting electrons $t_{\mathrm{cool}}$, and their gyroradius $r_{\mathrm{g,e}}$. There are 5 independent equations for these quantities; they are given in Sec. (2). Thus, there exist pairwise relations between the aforementioned parameters.
The most interesting is the relation between the gyroradius of plasma particles and typical scale of the magnetic field fluctuations. If $r_{\mathrm{g,pl}}$ turns out to be larger than $\lambda_{c}$, then the current-bearing particles freely move between fluctuations and the magnetic turbulence is linear. Otherwise, in the case $r_{\mathrm{g,pl}}<\lambda_{c}$, most of the plasma particles are trapped in regions of stronger magnetic field and the turbulence is non-linear. Linear turbulence can result from any instability (or combination of instabilities), both kinetic and MHD. Conversely, non-linear turbulence ultimately points at some MHD mechanism, whereas kinetic instabilities can play only auxiliary role. As we argue in this paper, rather general arguments lead to conclusion that the magnetic field turbulence in GRB afterglow blast waves has non-linear character. This has serious implications for understanding physics of relativistic shocks in general, and for GRB afterglow theories in particular.
2 Relations between shock’s microscopic parameters
Efficient synchrotron emission implies that the magnetic field does not decay faster than electrons lose their energy radiatively. So, the source must satisfy
$t_{\mathrm{dec}}\gtrsim t_{\mathrm{cool}}\ .$ (1)
Decay time for the turbulent magnetic field depends on the typical wavenumber of the fluctuations (i.e. on their inverse correlation length $\lambda_{c}$) and can be estimated from dispersion relation for Weibel instability modes assuming zero anisotropy (Morse & Nielson, 1971) [^1] We take Eq. (6) from this paper in the limit, where $\omega/(k_{y}v_{y})\ll 1$, thermal velocity is $v_{y}=c$, and ignore numerical factor $\pi^{-1/2}$. Note that $k_{y}$ in their notation is equivalent to our $\lambda_{c}^{-1}$. Since the solution of their dispersion relation is aperiodic, their $\omega$ is the same as $t_{\mathrm{dec}}^{-1}$.:
$t_{\mathrm{dec}}\simeq\frac{\omega_{\mathrm{pl}}^{2}\lambda_{c}^{3}}{c^{3}}\ .$ (2)
This relation may break for non-linear turbulence, and it is not clear beforehand whether the decay of non-linear fluctuations will be faster or slower.
The cooling time for synchrotron-radiating electrons is
$t_{\mathrm{cool}}\simeq\frac{6\pi\ m_{e}c}{\gamma\sigma_{\mathrm{{}_{T}}}B^{2}}=\frac{9}{4}\frac{m_{e}c^{3}}{\gamma e^{2}\omega_{\mathrm{{}_{B}}}^{2}}\ ,$ (3)
where $B$ is the magnetic field strength, $\gamma$ the Lorenz factor of emitting electrons, $\sigma_{\mathrm{{}_{T}}}$ Thomson cross-section, $m_{e}$ electron mass, $e$ elementary charge, and $c$ speed of light. The cyclotron frequency
$\omega_{\mathrm{{}_{B}}}\simeq\frac{eB}{m_{e}c}$ (4)
in turn can be expressed in terms of relativistic plasma frequency
$\omega_{\mathrm{pl}}\simeq\left(\frac{4\pi e^{2}Nc^{2}}{E_{\mathrm{pl}}}\right)^{1/2},$ (5)
where $N$ is number density of plasma particles and $E_{\mathrm{pl}}$ their energy. The ratio of these frequencies is
$\frac{\omega_{\mathrm{{}_{B}}}}{\omega_{\mathrm{pl}}}\simeq\frac{E_{\mathrm{pl}}}{m_{e}c^{2}}\left(\frac{B^{2}/4\pi}{NE_{\mathrm{pl}}}\right)^{1/2}=\sqrt{2\epsilon_{\mathrm{{}_{B}}}}\ \frac{E_{\mathrm{pl}}}{m_{e}c^{2}}\ ,$ (6)
where $\epsilon_{\mathrm{{}_{B}}}$ is the ratio of the magnetic-field energy density to plasma energy density.
Both emitting electrons and plasma particles are relativistic, so that their giroradii are given by similar expressions:
$r_{\mathrm{g,e}}=\gamma\ \frac{c}{\omega_{\mathrm{{}_{B}}}}\ ,\quad r_{\mathrm{g,pl}}=\frac{E_{\mathrm{pl}}}{m_{e}c^{2}}\ \frac{c}{\omega_{\mathrm{{}_{B}}}}\ ,$ (7)
and their ratio is
$\frac{r_{\mathrm{g,pl}}}{r_{\mathrm{g,e}}}=\frac{E_{\mathrm{pl}}}{\gamma m_{e}c^{2}}\ .$ (8)
Equations (1), (2), (3), (6), and (8) are the 5 independent equations relating the 6 unknown quantities of the problem. In addition, it is convenient to express the product $(\omega_{\mathrm{{}_{B}}}/\gamma)t_{\mathrm{cool}}$ using observable quantities:
$(\omega_{\mathrm{{}_{B}}}/\gamma)t_{\mathrm{cool}}\simeq\frac{9}{4}\frac{m_{e}c^{3}}{\gamma^{2}e^{2}\omega_{\mathrm{{}_{B}}}}=\frac{9}{4}\frac{\Gamma m_{e}c^{2}}{\alpha_{f}E_{\mathrm{sy}}}\ ,$ (9)
where $\alpha_{f}$ is the fine-structure constant and
$E_{\mathrm{sy}}\simeq\Gamma\gamma^{2}\hbar\omega_{\mathrm{{}_{B}}}$ (10)
the observed energy of photons in the synchrotron peak. We also consider the shock’s $\Gamma$ to be an observable, as it allows determination from spectrum and lightcurve modelling in a fairly independent way.
Now we re-write the efficiency condition (1) in terms of $(r_{\mathrm{g,pl}}/\lambda_{c})$ ratio using Eq. (2) to replace $t_{\mathrm{dec}}$:
$\frac{r_{\mathrm{g,pl}}^{3}}{\lambda_{c}^{3}}\lesssim\frac{\omega_{\mathrm{pl}}^{2}\ r_{\mathrm{g,pl}}^{3}}{c^{3}\ t_{\mathrm{cool}}}=\frac{\gamma^{3}\omega_{\mathrm{pl}}^{2}}{\omega_{\mathrm{{}_{B}}}^{3}t_{\mathrm{cool}}}\left(\frac{r_{\mathrm{g,pl}}}{r_{\mathrm{g,e}}}\right)^{3}\ .$ (11)
Substituting $(\omega_{\mathrm{{}_{B}}}/\omega_{\mathrm{pl}})$ from Eq. (6), $(r_{\mathrm{g,pl}}/r_{\mathrm{g,e}})$ from Eq. (8), and $(\omega_{\mathrm{{}_{B}}}/\gamma)t_{\mathrm{cool}}$ from Eq. (9) we arrive at
$\frac{r_{\mathrm{g,pl}}}{\lambda_{c}}\lesssim\left(\frac{2}{9}\frac{1}{\epsilon_{\mathrm{{}_{B}}}}\frac{\alpha_{f}E_{\mathrm{sy}}}{\Gamma m_{e}c^{2}}\right)^{1/3}\left(\frac{E_{\mathrm{pl}}}{\gamma m_{e}c^{2}}\right)^{1/3}.$ (12)
We conclude that plasma particles are confined by the magnetic field fluctuations and the magnetic turbulence in GRB afterglow emitting regions is essentially non-linear. Indeed, GRB afterglows typically have peak of their synchrotron emission in the few keV range (up to tens of keV in early afterglows), the only available determination of the magnetization parameter from the complete spectral analysis is $\epsilon_{\mathrm{{}_{B}}}\simeq 0.004$ (Asano et al., 2020; Derishev & Piran, 2021), and the numerical simulations tend to produce $\epsilon_{\mathrm{{}_{B}}}\sim 10^{-3}\div 10^{-2}$ (e.g. Chang et al., 2008; Sironi et al., 2015). Altogether, the first factor in the above inequality evaluates to $\sim 0.05$. The second factor cannot revert conclusion about smallness of $(r_{\mathrm{g,pl}}/\lambda_{c})$ ratio — this would require unreasonably low energy of emitting electrons, at least four orders of magnitude lower that the energy of plasma particles (a standard expectation is $(\gamma m_{e}c^{2})/E_{\mathrm{pl}}\gtrsim 1$).
3 Implications
There is common agreement that relativistic shocks propagating into unmagnetized medium (as in the case of GRB afterglows) generate their own magnetic field through Weibel-type (filamentation) instability, as suggested by Medvedev & Loeb (1999). The actual properties of the magnetic turbulence, such as the spatial scale of fluctuations and the level of magnetization, depend on what is the source of driving anisotropy in particles’ distribution. For example, predictions of the common collisionless shock scenario largely differ from predictions of the pair balance model, as discussed in the introduction. More important, however, is the fact that Weibel-type instability always saturates before entering non-linear phase (Garasev & Derishev, 2021) and therefore cannot explain non-linear turbulence required for GRB afterglows. This means that our understanding of processes that lead to generation of the magnetic field in relativistic shocks is essentially incomplete. Whatever is the outcome of Weibel-type instability, it should be followed by some kind of MHD instability capable of reaching the required level of non-linearity. And it is the latter that ultimately determines the level of magnetization. It is too early to speculate about the nature of this MHD instability.
In the pair-balance model, the magnetic field builds up due to continuous injection of electron-positron pairs, which occurs on the timescale approximately equal to $t_{\mathrm{cool}}$. A natural expectation is that the magnetic-field decay time is comparable to the build-up time, suggesting order-of-magnitude equality $t_{\mathrm{cool}}\sim t_{\mathrm{dec}}$ instead of relation (1). If so, then Eq. (12) should be treated as an estimate for the $(r_{\mathrm{g,pl}}/\lambda_{c})$ ratio rather than the upper limit for it.
Generally speaking, one should not treat our result as a proof that the magnetic-field turbulence is strongly non-linear. It is possible that the actual fluctuations are just non-linear enough to form soliton-like structures, which could persist much longer than linear perturbations of the same scale.
Another important question is whether the magnetic turbulence is linear with respect to the emitting electrons. As follows directly from Eq. (12), via substitution of $(r_{\mathrm{g,pl}}/r_{\mathrm{g,e}})$ from Eq. (8),
$\frac{r_{\mathrm{g,e}}}{\lambda_{c}}\lesssim\left(\frac{2}{9}\frac{1}{\epsilon_{\mathrm{{}_{B}}}}\frac{\alpha_{f}E_{\mathrm{sy}}}{\Gamma m_{e}c^{2}}\right)^{1/3}\left(\frac{\gamma m_{e}c^{2}}{E_{\mathrm{pl}}}\right)^{2/3}.$ (13)
This indicates that for the emitting electrons the turbulence is still non-linear at least at the early afterglow stage. The pair-balance model predicts increase of $\gamma$ with time, roughly following $B^{-1/3}$ law, while $E_{\mathrm{pl}}$ is expected to decrease proportionally to the shock’s Lorentz factor $\Gamma$. Then, for the late stages of afterglow evolution, one may expect that the turbulence becomes linear with respect to the emitting electrons (being non-linear with respect to the plasma particles).
In any case, the emitting electrons are loosely confined in the magnetic field inhomogeneities. Electrons’ displacement due to diffusion can be estimated as
$\ell_{\mathrm{diff}}\simeq\sqrt{Dt_{\mathrm{cool}}}\simeq r_{\mathrm{g,e}}\sqrt{\omega_{\mathrm{{}_{B}}}t_{\mathrm{cool}}/(3\gamma)}\ ,$ (14)
where we assumed Bohm diffusion coefficient $D=cr_{\mathrm{g,e}}/3$. Substituting $r_{\mathrm{g,e}}$ from Eq. (13) and $(\omega_{\mathrm{{}_{B}}}/\gamma)t_{\mathrm{cool}}$ from Eq. (9), we find that
$\frac{\ell_{\mathrm{diff}}}{\lambda_{c}}\lesssim\left(\frac{1}{48}\frac{1}{\epsilon_{\mathrm{{}_{B}}}^{2}}\frac{\Gamma m_{e}c^{2}}{\alpha_{f}E_{\mathrm{sy}}}\right)^{1/6}\left(\frac{\gamma m_{e}c^{2}}{E_{\mathrm{pl}}}\right)^{2/3}\sim 30\left(\frac{\gamma m_{e}c^{2}}{E_{\mathrm{pl}}}\right)^{2/3}.$ (15)
Here we take $\epsilon_{\mathrm{{}_{B}}}=0.004$, $E_{\mathrm{sy}}=5$ keV, and $\Gamma=30$ to estimate the first factor. Thus, the emitting electrons travel diffusively far enough to sample many irregularities of the magnetic field over their cooling time. This means that one-zone approximation for GRB afterglow emitting region can still be used despite non-linear character of the magnetic turbulence. However, it is necessary to average the synchrotron spectrum of an individual electron over regions with different magnetic field strength. It is hard to guess a priori what could be the distribution function for local values of the magnetic field strengths, and this distribution may strongly deviate from Gaussian, expected for a linear turbulence.
Acknowledgements
This work was supported by the Russian Science Foundation under grant no. 24-12-00457.